If 1.6 accidents can be expected an intersection on any given day, what is the


probability that there will be 3 accidents on any given day?

1
Expert's answer
2022-04-14T00:58:41-0400

We have a Poisson distribution,

"λ=1.6;\\
P_t(X=k)=\cfrac{(\lambda t)^k\cdot e^{-\lambda t}}{k!}=\cfrac{(1.6t)^k\cdot e^{-1.6t}}{k!},\\
P_1(X=3)=\cfrac{(1.6\cdot1)^3\cdot e^{-1.6\cdot1}}{3!}=0.1378."